Find the quadratic model for the set of values.
(1,-2), (2,-2), (3,-4)

Answer by stanbon(48558)

Find the quadratic model for the set of values.
(1,-2), (2,-2), (3,-4)
----------------------------------
Substitute each of those x/y pairs into the
quadratic form y = ax^2+bx+c.
You will get three equations with variables a,b, and c.
Solve the system of equations.
Rewrite y = ax^2+bx+c with the values you find for a,b,and c.
Cheers,
Stan H.
Answer: y = -x^2+3x-4
=====================